Home‎ > ‎Crew 162‎ > ‎

0202-DailySummaryReport

DAILY SUMMARY REPORT SOL 10

Summary Title: Clear skies after the storm

Mission Status: Hab is operational, systems are in a good config. Mission
objectives are on track for a successful second week.

Sol Activity Summary: Our crew excitedly welcomed our visiting crew member,
Robert, a professional photographer from the UK. His spaceship landed just
before 10:00 and he joined our morning daily planning conference. Our EVA
yielded helpful navigational data and some incredible shots of our team
hiking across the snow-covered Martian hills near the habitat. The skies on
Mars are clear tonight and we will be observing with the telescope. We got
some great photos of the ISS flying over the hab before the opening of
comms.

Look Ahead Plan: We are nearly half-way through our second week...time
flies on Mars! We have only 4.5 days left on in our mission.

Anomalies in work: No anomalies to report.

Weather: Snow has stopped and skies have cleared. Sunny and cold day, cold
night with light wind. No clouds.

Crew Physical Status: One crewmember reporting shoulder pain from the EVA
backpack not being adjusted correctly; we will switch her pack for
tomorrow's EVA. Otherwise crew is healthy and happy.

EVA: EVA #4 was a success. Summary report will be submitted tonight. EVA #5
request for Sol 11 will be sent tonight also.

Reports to be filed:
1. Commander Report
2. Engineering Report
3. Journalist Report
4. Daily Photos
5. Greenhab Report
6. EVA #4 Summary Report
7. EVA #5 Request
No astronomy report but we will be firing up the telescope tonight. Report
to follow tomorrow.

Support Requested: Help in locating mouse traps but other than that nothing
at this time